OpenStudy (anonymous):

Please help! A particular wire has a resistivity of 5.63×10-8 Ωm and a cross-sectional area of 2.16 mm2. A length of this wire is to be used as a resistor that will develop 130 W of power when connected to a 7.50 V battery. What length of wire is required?

OpenStudy (anonymous):

Do you know that power \(P = V^2/R\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I found R to equal .433ohms

OpenStudy (anonymous):

and then (.433)(2.16E-6)/5.63E-8 and this equals 1.66E-15

OpenStudy (anonymous):

but it keeps telling me its wrong

OpenStudy (anonymous):

Your approach is correct! But your calculation is wrong. I think you have added the power of 10. I am getting 16.61 meter.

OpenStudy (anonymous):

Oh thank you! so, where was I adding the power of ten wrong?

OpenStudy (anonymous):

You were doing -6 + (-8) = -14 but actually it should be -6-(-8) = 2

OpenStudy (anonymous):

oh, thank you!

OpenStudy (anonymous):

could you help me with my next question? I posted it if you want the best response again?